OpenStudy (anonymous):

How do you integrate this please?? dr/dt = 2r(1-r)

OpenStudy (anonymous):

nice job mimi

OpenStudy (lalaly):

\[\frac{dr}{2r(1-r)} = dt\]\[\int\limits{\frac{dr}{2r(1-r)}} = \int\limits{dt}\]

OpenStudy (anonymous):

[1/2r(1-r)]dr=dt =>integrating on both sides 1/2log r -1/2 log (1-r)=t

OpenStudy (lalaly):

the first part by partial fractions \[\int\limits{dt} = t\]

OpenStudy (anonymous):

Thanks guys!